At what altitude above the Earth's surface is the acceleration due to gravity equal to g/132?
Answer #1

9.81 / 132 = (G * 5.9 * 1024) / r2

9.81 / 132 = (6.67 * 10-11 * 5.9 * 1024) / r2

r = 7.27 * 107 m

The radius of earth = 6.3675 * 106

altitude = (7.27 * 107) - (6.3675 * 106 )

= 6.64 * 107 m